Status Report

ISS On-Orbit Status 23 Dec 2002

By SpaceRef Editor
December 23, 2002
Filed under , ,
ISS On-Orbit Status 23 Dec 2002

All ISS systems continue to function nominally, except as noted previously or below.  Start of Week 4 for Expedition 6.

After wake-up at the regular 1:00am, the crew was commended on keeping busy yesterday, a rest day (perhaps anticipating Santa?).

FE-2/SO Don Pettit was scheduled for on-orbit calibration of the RED (resistive exercise device) Flexpack canister settings.

FE-1 Nikolai Budarin spent part of the day working in the DC-1 “Pirs” docking compartment, first cleaning the VD1 and VD2 air ducts to the Service Module (SM), later performing the regular (monthly) check-up of the circuit breakers (AZS) on the BVP amp switch panel — they should all be On — and the 14 LEDs of the fuses in fuse panels BPP-30 and BPP-36.  He also changed out the filters of the DC-1 PF1 and PF2 dust collectors, followed by clean-up of the V1 and V2 fans.

Budarin, assisted by Pettit as CMO (crew medical officer) was scheduled to complete a session with the MO-5 MedOps protocol of cardiovascular evaluation during graded exercises on cycle ergometer.  [Using the Gamma-1 ECG equipment with biomed harness, skin electrodes, and a blood pressure and rheoplethysmograph cuff wired to the cycle ergometer’s instrumentation panels, the cosmonaut was to work the pedals after a prescribed program at load settings of 125, 150, and 175 watts for three minutes each.  All measurements were to be recorded and telemetered to MCC-M, from where the workout was controlled.]

CDR Ken Bowersox started on the FOOT (Foot/Ground Reaction Forces During Space Flight) experiment, first by preparing himself with a CBT (computer-based training) session, then setting up the hardware and preparing for EMG (electromyography, i.e., muscular electrical signals) calibration.  Today‚s activities got the hardware ready for tomorrow‚s FOOT activities.

FE-1 Budarin was to perform hardware checkout for the Lada-2 (Plants-2) experiment, dismantling, filling and reinstalling the KDV water container, then watering the substrate with tensiometer monitoring of the substrate humidification, and preparing data for downlinking. Plants-2 will be continued tomorrow, with plant seeding and configuring of the payload for cultivation.

Budarin also was scheduled to conduct the periodic inspection of the SM pressure hull in the Working Compartment behind panels 130, 134, 135, 138, and TVIS for moisture, mold, or corrosion.  [Nikolai was to use surface wipes to clean the area in question, describe any discovered mold as to color, consistency (solid, crumbly, powdery), and surface area, and photograph the hull surface with the digital camera before and after removing the mold.  Pictures were to be downlinked to MCC-H to be forwarded to MCC-M.]

The crew performed their daily physical exercise program.

The recent low-capacity trip of SM battery 5 was expected due to lifetime limits.  But the trip of SM battery 7 trip was unexpected; since specialists had calculated approximately ten more months of lifetime left.  Now they believe battery 7 is hard failed.  MCC-M is proposing to change-out battery 5 in the near future, but to defer the battery 7 change-out until Progress 10P arrival.

After the recent successful ARCTIC 1 refrigerator/freezer screw extraction and troubleshooting procedure by the Don Pettit, review of the downlinked data indicates that a thermo-electric unit is failed. There may be no additional on-orbit troubleshooting options at this time.

Correction of yesterday’s Status Report:  Saturday’s earlier-than-expected maneuver to YVV/”barbecue” maneuver, surprisingly enough, was caused by temperature exceedance not on the Progress batteries, but on the ISS truss hardware.  [MCC-H and MCC-M had anticipated that the Progress batteries would register overtemperatures, but their limit was not quite reached yet when truss hardware, such as S-band, surprisingly started to exceed.  As of now, the “sidewise” YVV attitude is working very well for the current solar Beta condition; and temperatures are nominal. However, YCC is costing us about 100 kg of prop now instead of the 30 kg originally calculated for constrained-by-Progress case due to the longer time spent in YCC on thrusters..  Another decidedly positive step on the learning-by-experience curve we have been on since Day One.]

Today’s targets for the CEO (crew earth observations program) were Eastern Mediterranean Dust and Smog (on this ascending pass, the crew was to look obliquely to the right of track and try to detect sunlit aerosol plumes over the darker waters of the Mediterranean), Western Mediterranean Dust and Smog (with a strong winter storm forecast to be approaching from the west, crew was to look right of track for plumes of dust borne on strong southerly winds of the Sahara and over the dark waters of the Med), Mount Etna (DYNAMIC EVENT SITE:  If the crew looked left of track, they had good views of the sunlit plume of this active volcano), Great Lakes in Glint (DYNAMIC EVENT SITE:  Looking obliquely to the right of track this pass for what should have been spectacular panoramic views of the Great Lakes in sun glint), and Chicago, Illinois (crew had a nadir view of the city in a wintry setting with ice and possible light snow cover).

SpaceRef staff editor.